OpenStudy (anonymous):

On a $12.00 purchase, Michael was offered 3 successive discounts of 25%, 20% and 10% each in any order he wished. What is the difference in dollars between applying the discounts in order from greatest to least and in order from least to greatest?

OpenStudy (lilg132):

12 / 100 = 0.12 (This is 1% of 12) 0.12 * 25 = 3 (this is 25% of 12) 12 - 3 = 9 (this is how much the price will be after 25% discount

OpenStudy (lilg132):

now to work out how much the price will be after 20% of 9 we do the same thing instead of using 12 we use 9

OpenStudy (lilg132):

9 / 100 = 0.09 (this is 1% of 9) 0.09 * 20 = 1.8 (this is 20% of 9) 9 - 1.8 = 7.2 (this is the price after 20% of 9)

OpenStudy (anonymous):

1.8?

OpenStudy (lilg132):

now to get 10% of that 7.2 / 100 = 0.072 (this is 1 percent) 0.072 * 10 = 0.72 (this is 10 percent) 7.2 - 0.72 = 6.48 this is the price after taking all discounts out now including 10% from the order of first 25% discount then 20% discount and then 10% discount

OpenStudy (lilg132):

now the question wants you to take off th discounts in the opposite order so first you need to take of 10% then 20% then 25% but remember start with the beginning price which is 12 will you manage to do this now? if not please tell me and i will go through it

OpenStudy (lilg132):

1.8 was 20% of $9

OpenStudy (anonymous):

no, not realy

OpenStudy (lilg132):

ok have you understood what i have done so far?

OpenStudy (anonymous):

yes...

OpenStudy (lilg132):

if you haven't please tell me

OpenStudy (lilg132):

ok lets do the next part of the question where it says do the discounts in reverse order which means take off 10% from the price then 20% and then 25%

OpenStudy (lilg132):

12 / 100 = 0.12 (this is 1%) to work out what 10% is of the price we multiply the 1% by 10 0.12 * 10 = 1.2 (this is 10% of the price $12) now to get the price after discount we minus the 10% from the price 12 - 1.2 = $10.8 (this is the price after 10% discount)

OpenStudy (lilg132):

now we need to take off 20% off the price 10.8 / 100 = 0.108 (this is 1%) 0.108 * 20 = 2.16 (this is 20%) 10.8 - 2.16 = 8.64 (this is the price after 20% discount)

OpenStudy (lilg132):

now we need to take off 25% discount from the price 8.64 / 100 = 0.0864 (1%) 0.0864 * 25 = 2.16 (25%) 8.64 - 2.16 = $6.48

OpenStudy (lilg132):

so now what was the difference in price if we did the discounts the first way or the way i did it now?

OpenStudy (anonymous):

0

OpenStudy (lilg132):

yes so there is no difference in which order you take the discounts

OpenStudy (anonymous):

I did this a more simplier way!

OpenStudy (lilg132):

i was doing it in a way you might understand i know you may have just done 12 * 0.25

OpenStudy (anonymous):

ya, i did that , that way! :)

OpenStudy (lilg132):

ok next time if you have any problems i will remember that you understand that way :)

OpenStudy (anonymous):

:) thanks so much!!!!!!

OpenStudy (lilg132):

your welcome